OpenStudy (anonymous):

5/2y=(y+5) how do u solve it??

OpenStudy (anonymous):

first step multiple both sides by 2y

OpenStudy (kkutie7):

\[\frac{ 5 }{ 2y }=(y+5)\] \[5=2y(y+5)\] \[5=2y ^{2}+10y\] \[0=2y^{2}+10y-5\] now factor

OpenStudy (anonymous):

how do u factor that?? help :( thanks!

OpenStudy (kkutie7):

This wont factor so you need to use the quadratic formula ok

OpenStudy (kkutie7):

A=2,B=10,C=-5 \[y=\frac{ -b \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

it would be 150/2 or -130/2 right?

OpenStudy (anonymous):

so whats the answer then?

OpenStudy (kkutie7):

its at the bottom i believe. i don't have a calculator right now or id solve it for you.